OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (campbell_st):

it just means to substitute n =1, n = k and n = k + 1 into the right hand side. \[P_{1} = 1\] \[P_{k} = k(k +1)\] \[P_{k + 1} = (k +1)(k+1+1)\] = (k + 1)(k +2) it looks a little bit like mathematical induction.